# Hierarchical Sequence Iteration for Heterogeneous Question Answering

### v1

@@ A.4 Notations
$\pi_{\theta}$ HSEQ-I iteration policy (LLM-based) with parameters $\theta$; operates over $(q,S_{h})$ to select evidence iteratively.
@@
$\mathcal{H}$ HSEQ-H head module for answer synthesis from $(q,\kappa(M_{\tau}))$; can also generate guidance $g$.
@@
A.5 Example Using HSEQ
A.5.1 Case Study: Guided Iterative Retrieval on HybridQA
@@
Query $q$: “Who is the author of the novel that inspired the 2004 Russian film directed by Timur Bekmambetov?” HSEQ-I (iterator): Qwen3-4B-Instruct-2507; HSEQ-H (head): Falcon-H1-7B-Instruct. Guidance mode: head; source: cache (latency $\approx$ 0.12 ms).
@@
A.5.2 Case Study: Guided Iterative Retrieval on HotpotQA
@@
Query $q$: “Which style is the building located on the East Side of Midtown Manhattan that Robert Von Ancken appraised?” HSEQ-I (iterator): Qwen3-4B-Instruct-2507; HSEQ-H (head): Falcon-H1-7B-Instruct. Guidance mode: head; source: generated online (latency $\approx$ 8,496 ms).
@@
A.6 Statement for The Use of Large Language Models (LLMs)
We used large language models (LLMs) as general-purpose tools for writing assistance and engineering support. For writing, we employed LLMs to improve clarity and style (e.g., rephrasing sentences, tightening paragraphs, standardizing notation, and proofreading grammar). Drafting, technical claims, algorithms, proofs, experiment design, and all final wording were authored and verified by the authors. For engineering, we consulted LLMs for debugging; all research code, data processing, and experiment scripts were implemented, audited, and executed by the authors. No text or code generated by an LLM was used verbatim without author review; we take full responsibility for the content.

### v3

@@ A.4 Notations
$\pi_{\theta}$ RELOOP-I iteration policy (LLM-based) with parameters $\theta$; operates over $(q,S_{h})$ to select evidence iteratively.
@@
$\mathcal{H}$ RELOOP-H head module for answer synthesis from $(q,\kappa(M_{\tau}))$; can also generate guidance $g$.
@@
A.5 Overall performance of RELOOP agent pairs
Table 8 shows the overall performance of RELOOP agent pairs on Hybrid-QA, containing Accuracy/F1 and Efficiency.
Table 8: Overall performance of RELOOP agent pairs on Hybrid-QA: Accuracy/F1 and Efficiency.
Accuracy & F1 Efficiency
Iteration Agent (RELOOP-I) Head Agent (RELOOP-H) Avg. Acc Avg. F1 Steps $\downarrow$ Latency (ms) $\downarrow$
Falcon3-3B-instruct Falcon3-10B-instruct 59.3 61.1 2.6 10530.1
Falcon3-3B-instruct Llama-3.1-8B-Instruct 61.2 65.1 2.46 11616.7
Falcon3-3B-instruct Falcon-H1-7B-Instruct 56.1 58.6 2.25 11714.4
A.6 Example Using RELOOP
A.6.1 Concrete fragment examples.
Below we instantiate $(\ell,p,c,\mu)$ for three modalities (all values are illustrative):
$s_{\text{text}}$ $=\big(\mathrm{id}=s_{1},\,\ell=\texttt{sentence},\,p=p_{1},\,c=\text{``...capital is Paris...''},$
$\qquad\mu=\{\,\begin{aligned} &\texttt{source\_id}=\texttt{doc\_12},\\ &\texttt{offsets}=(128,172),\ \texttt{time}=\texttt{1992}\}\end{aligned}$
$s_{\text{row}}$ $=\big(\mathrm{id}=r_{3},\,\ell=\texttt{table\_row},\,p=t_{1},\,c=\text{``France | 67.4M | EU''},$
$\qquad\mu=\{\,\begin{aligned} &\texttt{source\_id}=\texttt{tbl\_7},\,\texttt{schema}=\text{[country,pop,bloc]},\\ &\texttt{row}=3\}\end{aligned}$
$s_{\text{cell}}$ $=\big(\mathrm{id}=u_{3,2},\,\ell=\texttt{table\_cell},\,p=r_{3},\,c=\text{``67.4M''},$
$\qquad\mu=\{\texttt{source\_id}=\texttt{tbl\_7},\,\texttt{row}=3,\,\texttt{col}=2\}\big);$
$s_{\text{kg}}$ $=\big(\mathrm{id}=k_{9},\,\ell=\texttt{triplet},\,p=g_{1},\,$
$\qquad c=\text{``(Paris, capital\_of, France)''},\mu=\{\texttt{source\_id}=\texttt{kg\_2}\}\big).$
Segments are concatenated in parent-before-child order. This minimal contract enables structure-aware neighborhoods and budget-aware iteration without inspecting raw files.
A.6.2 Case Study: Guided Iterative Retrieval on HybridQA
@@
Query $q$: “Who is the author of the novel that inspired the 2004 Russian film directed by Timur Bekmambetov?” RELOOP-I (iterator): Qwen3-4B-Instruct-2507; RELOOP-H (head): Falcon-H1-7B-Instruct. Guidance mode: head; source: cache (latency $\approx$ 0.12 ms).
@@
A.6.3 Case Study: Guided Iterative Retrieval on HotpotQA
@@
Query $q$: “Which style is the building located on the East Side of Midtown Manhattan that Robert Von Ancken appraised?” RELOOP-I (iterator): Qwen3-4B-Instruct-2507; RELOOP-H (head): Falcon-H1-7B-Instruct. Guidance mode: head; source: generated online (latency $\approx$ 8,496 ms).
@@
A.7 Statement for The Use of Large Language Models (LLMs)
We used large language models (LLMs) as general-purpose tools for writing assistance and engineering support. No text or code generated by an LLM was used verbatim without author review; we take full responsibility for the content.